A charge of 6 ร— 10โˆ’7๐ถ is transferred from infinity to point B. If the work




done by the electric force to do this is 1.2 ร— 10โˆ’5๐ฝ. What is the potential at point




B?

1
Expert's answer
2022-03-07T12:21:34-0500

Answer

the potential at point

"V=\frac{W}{q}\\=\frac{1.2*10^{-5}}{6*10^{-7}}\\=20Volts"